OpenStudy (newarf):

y^5*y^14= 2y^16 y^60 y^19 2y ^60

OpenStudy (jiteshmeghwal9):

use the property\[a^b.a^c=a^{b+c}\]

OpenStudy (newarf):

so y ^19

OpenStudy (jiteshmeghwal9):

yes

OpenStudy (newarf):

ok I have another one can you help

OpenStudy (phi):

you use the same idea as y* (y*y) = y*y*y \[ y^1 \cdot y^2 = y^3\]

OpenStudy (newarf):

my next question is w^12/ w^18

OpenStudy (jiteshmeghwal9):

use the property \[\frac{a^b}{a^c}=a^{b-c}\]

OpenStudy (newarf):

ok thanks